The sum of 4 hours and 17 minutes, 3 hours and 58 minutes, 45 minutes, and 7 hours and 12 minutes is?
Answer by Rollerprincess(4)

1. total the whole hours. 4 + 3 + 7 = 14
2. total the minutes. 17 + 58 + 45 + 12 = 132
3. convert 132 minutes into hours - 132/60 = 2 hours and 12 minutes.
4. Total the hours from #1 and the hours from #3. 14h + 2h = 12m = 16 h 12 m.